The perimeter of a basketball court is 102 meters. The length of the court is 6 meters more than twice the width. What are the dimensions of the court?
Answer by nerdybill(7384)

The perimeter of a basketball court is 102 meters. The length of the court is 6 meters more than twice the width. What are the dimensions of the court?
.
Let w = width of court
then
2w+6 = length of court
.
2(w+ 2w+6) = 102
2w+ 4w+12 = 102
6w+12 = 102
6w = 90
w = 15 meters (width)
.
Length:
2w+6 = 2(15)+6 = 30+6 = 36 meters (length)
.
dimensions:
15x36 meters